A cable stretches by an amount x under a certain load. if it is replaced by a cable of the same material but half as long and half the diameter, the same load will stretch it by ?
1
Expert's answer
2019-04-08T09:45:56-0400

Young's modulus is


"E=\frac{FL}{Ax}=\frac{0.5LF}{A'x'}"

"\frac{FL}{\pi r^2x}=\frac{0.5LF}{\pi (0.5r)^2x'}"

"x'=2x"